OpenStudy (anonymous):

ABCD is a quadrilateral with A(8,-1), B(10,-7), C(-6,-1), and D(0,2). Determine whether ABCD is a trapezoid.

OpenStudy (anonymous):

Find the slope of both sides.... thanks open study team!

OpenStudy (anonymous):

yes it is a trapezoid

OpenStudy (anonymous):

which sides

OpenStudy (anonymous):

sides AC and BD?

OpenStudy (anonymous):

the slope for AC= 9/5 BD= 6/1

OpenStudy (anonymous):

So its not a trapozoid?

OpenStudy (anonymous):

it is a trapezoid

OpenStudy (anonymous):

OOOHHHHH